OpenStudy (anonymous):

how much of an alloy that is 40% copper should be mixed with 600 ounces of an alloy that is 70% copper in order to get an alloy that is 50% copper? ans in ounces pls

OpenStudy (anonymous):

Y oz alloy1 gives .40(Y) oz Cu 600 oz alloy2 gives .70(600)=420 oz Cu Total Cu = 420+.40Y Total weight = 600+Y %Cu =( 420+.40Y)/(600+Y) = .50 .5(600+Y)=420+.40Y .1Y=120 Y=1200 oz